### v1

@@ Abstract
The process of developing theories and models and testing them with experiments is fundamental to the scientific method. Automating the entire scientific method then requires not only automation of the induction of theories from data, but also experimentation from design to implementation. This is the idea behind a robot scientist—a coupled system of AI and laboratory robotics that has agency to test hypotheses with real-world experiments.
In this chapter we explore some of the fundamentals of robot scientists in the philosophy of science. We also map the activities of a robot scientist to machine learning paradigms, and argue that the scientific method shares an analogy with active learning.
We demonstrate these concepts using examples from previous robot scientists, and also from Genesis: a next generation robot scientist designed for research in systems biology, comprising a micro-fluidic system with 1000 computer-controlled micro-bioreactors and interpretable models based in controlled vocabularies and logic.
Keywords:
Robot scientist scientific discovery active learning laboratory robotics.
@@ 1 Introduction
The use of AI-robotic systems in scientific research has been demonstrated not only possible, but fruitful, in the past two decades, with several projects—from Adam [16] and Eve [34] to the Robot Chemist [6]—proving the value of coupling AI software agents with experimental platforms to give them real-world agency. Research toward the goal of productive robot scientists is necessarily a multi-disciplinary endeavour. Fields that contribute include: artificial intelligence, robotics, nanotechnology and materials science. For robot scientist projects to be a success, researchers need a shared understanding of the process they are working to automate: the scientific method.
We begin this chapter by defining some of the core concepts in the philosophical ideas behind robot scientists in Section 2, beginning with theories and models, the fundamental entities in scientific research. We explore the concept of parsimony as it relates to the scientific method, something which is learned through scientific education in humans but requires explicit care when creating robot scientists. Then we cover methods of inference, and the difference scientific values which we use to assess the value of a theory or model. Section 3 aims to identify which machine learning paradigm is most apt for scientific discovery. Ultimately, we conclude that scientific discovery shares an analogy with active learning, and that this, rather than reinforcement learning, is the most useful paradigm to adopt when designing or analysing robot scientists. Section 4 gives an overview of systems biology, the domain in which the next generation robot scientist Genesis is applied. We argue that as biological systems are “complex systems” as described by [33], they are excellent targets for study by robot scientists with their superhuman abilities in reasoning and precision. In Section 5 we present a case study of Genesis, and one model that we have developed with the aim of automated scientific discovery.
@@ 2 The Philosophy of Robot Scientists
A robot scientist is defined in [16] as
a physically implemented laboratory automation system that exploits techniques from the field of artificial intelligence to execute cycles of scientific experimentation.
@@
The goal of science is to develop theories that explain and predict phenomena in the real world. To develop theories science uses models, which are representations of theories in some localised context. Models are a surrogate for the system being studied (object system), in the sense that they display characteristics or behaviour, of sufficient similarity to that of the object system, that allow for indirect study of the object system by studying its surrogate (the subject system). Models are particularly useful when direct study of a system is impossible, impractical or undesirable.
@@
In both of these cases, the object system is the same: a cell. However they have quite different qualities. Model 1 would be well-suited to teaching high school students how cells of, for example, the yeast Saccharomyces cerevisiae function. However, Model 2 is capable of quantitative predictions, enabling direct comparison with quantitative experimental data.
@@
Models with deductive capacity using mathematics, such as Model 2, are useful for all forms of scientific discovery, but particularly when automated. The scientific discovery problem becomes
@@
an incremental process of refinement [of the model] strongly guided by the empirical observations [9].
@@
Methods of scientific enquiry rely on: constructing a good starting model; inferring changes to the model; techniques to reason about which models are better; and the collection of relevant, high-quality empirical data to drive the process.
@@
The purpose of a robot scientist is to provide software and hardware that together can achieve each each of these activities, and join them together in a “closed-loop” form of enquiry without human intervention [16]. Through the process of designing a system capable of independent scientific inquiry, we seek insight into the scientific method itself, as well as the system subject to enquiry.
2.1 Core aspects of scientific method
@@
There are three components of the scientific method that we will consider here, namely: parsimony, logical inference, and statistical inference. A history and detailed treatment of each of these components is covered in [10].
@@ 2.1.1 Parsimony
Parsimony is the concept that when choosing between theories that fit the data equally well, the theory that is simplest is preferable. Sometimes referred to as “Ockham’s razor”, parsimony is absolutely essential for science, and it is easy to take this principle as a common sense notion and end further rumination on the subject. However, a proper treatment of the motivations for adopting parsimony as a guiding principle is warranted when it comes to the automation of discovery and systems biology.
Firstly, it is worth distinguishing epistemological parsimony—which we alluded to above, that between theories that fit data equally well the simplest is preferable—from ontological parsimony—that nature itself prefers simplicity. Arguments for ontological parsimony suffer from many counter-examples, for example the yeast genome went through a duplication during its evolution resulting in numerous genes with overlapping or identical function [15]. However all scientific argument must adopt some version of ontological parsimony, as it is the basis for generalisation of theories.
@@
Concrete examples of ontological parsimony include: taking the assumption that there exist common properties between individuals of the same species. This could be a chemical species, that all glucose molecules react with water molecules in the same ways. Or it could be a species of organism, implying that we can apply the same computational model to two individual colonies of S. cerevisiae.
These ideas of ontological parsimony are so central to scientific enquiry that researchers may not consider them—they become necessary implicit biases.
@@
One power of appealing to ontological parsimony is that it is the basis of factorial experimentation, described in [7] as a two-step process: first we identify relevant factors for the phenomenon to be studied; then we design experiments holding certain factors constant and varying over others. Determining relevant factors means stating that certain factors are irrelevant, that we do not expect them to affect the outcome of the experiment. Deciding which factors to include is not an easy process. Designing experiments in this way allows the controlled study of phenomena so that we can test hypotheses that relate to a restricted subset of factors. Factorial design of experiments requires an appeal to ontological parsimony, informally
@@
we expect empirical data from experiments of the same class to exhibit only random variation,
@@
where the class of experiment depends only on the variable factors. This assumption allows the inference of empirical laws about the phenomenon.
@@
In practice we cannot usually keep constant all the factors we would like, so the class of experiment does not only depend on relevant variables. The nature of a given experimental protocol will introduce systematic errors: variations in the empirical data not arising from the experimental variables or random noise. Many methods exist to mitigate and model systematic errors, using randomisation techniques, systematic design, and statistics. Modelling systematic errors allows us to create a map on to each class of experiments as defined above. Randomisation, or systematic designs like Latin squares, can be impractical when using robot scientists due to the limitations of the automation hardware. One example from biology is that liquid handling procedures for a Latin square are impossible on certain liquid handling robots, and can increase procedure times by orders of magnitude on those with such a capability.
@@
On the other hand, robot scientists have advantages when it comes to relying on and examining this underlying bias. Firstly, that robots are capable of performing repeated tasks with a much higher accuracy and precision than human counterparts, as is shown by examples of adoption of laboratory automation in biology [12], physics [28], and chemistry [32]. Secondly, that we can evaluate the validity of this assumption by recording more data about the execution of experiments than is usually recorded when humans complete experiments. This is a “natural by-product” [16], as robots and software frequently have automatic logging capability. Finally, that we are forced to encode the experimental protocol forces the scientist to specify which experimental factors will be constant, or that if they are not constant they are judged irrelevant, and therefore they do not break the validity of statement above by creating a new class of experiments (for example, the use of different individual glass flasks of the same brand, model, and age).
@@
Epistemological parsimony is a concept much more familiar to human scientists, or at least one which is applied more explicitly. It has been formalised in information theory as the minimum message length (MML) by combining Shannon’s formula for the information ($I$) of an event ($E$), given by $I(E)=-log_{2}(P(E))$, with Bayes’ theorem, $P(E_{1}\cap E_{2})=P(E_{1})P(E_{2}|E_{1})$. MML states that the hypothesis $H$ that best explains data $D$—in other words it maximises $P(H\cap D)$—is that which minimises the information (message length): $I(H\cap D)=I(H)+I(D|H)$. In this form we can see there is a trade-off between the first and second term meaning information is only added to a hypothesis if we can use this information to explain the data, which is the principle of Ockham’s razor [2, 31, 1].
@@
Epistemological parsimony is the basis for several fundamental concepts in the modern scientific method. One example is the idea of a null hypothesis—we present a null hypothesis stating that any variance in empirical data can be explained by our extant model, to a degree of certainty due to random noise. To reject the null hypothesis is to accede a more complex model, and we only do so if we are certain our current model is insufficient to explain the empirical data.
2.1.2 Logical inference
Formally, logics are mathematical languages that relate premises and conclusions. There are various logics that can be used for reasoning in science, including propositional logic and first-order logic. We use logics in science to represent facts (observations of the real world) and laws (parts of theories or models) [7, 10].
There are three basic forms of logical inference. Firstly, deductive logic, where conclusions are derived from premises and laws. A valid deductive argument is one that guarantees the truth of its conclusions given the truth of its premises. Deductive logic is what enables deterministic simulations, and how we reason about a hypothesis in the scientific method.
Inductive logic and abductive logic cannot provide such guarantees. In both cases, we seek explanations for a set of observable facts. In the case of induction, we are seeking laws that can explain a general case, whereas in the case of abduction we are seeking facts that can explain specific cases. Both induction and abduction deal in uncertainty, and for that we need probabilistic and statistical inference.
2.1.3 Statistical inference and probability
Certain laws in science may only be expressed to a degree of certainty. For such situations we require statements that can deal with uncertainty. Carnap defines two types of probability [7]. Statistical probability is a frequentist idea about the relative frequencies of mass events. The defining characteristic of statements of statistical probability is that they cannot be decided by logic, but rest on empirical observations. Logical probability on the other hand is the probability on a logical relation between two propositions. Gauch, Jr prefers to describe the two types of probability as being about events and beliefs respectively [10].
Using statistics, we can reason to a model given observations. In the case of abduction, we are reasoning about facts; with induction, laws. The statistical element of this reasoning represents the uncertainty around the model, which can be viewed either as a belief in the law, or perhaps more commonly in empirical science, the relative frequencies of events. Either way, these processes is crucial to the scientific method because they allow us to form and improve upon models.
@@ 2.2 Comparing scientific models
The relative merit of competing scientific models is not a trivial assessment. This problem is referred to in the philosophy of science as the problem of theory choice. For a robot scientist to be effective we must incorporate values and an evaluation procedure, otherwise the scientific discovery process will depend on human evaluation.
@@
  • Internal consistency: defined as the absence of contradictions within a theory. This can be extended to include the various contexts in which the theory is applied; in biology this could mean growth of S. cerevisiae in different conditions.
  • External consistency: then the absence of contradictions with other scientific theories. For example, that our model of the biochemistry of yeast is thermodynamically consistent.
  • Empirical accuracy: otherwise referred to as predictive power, this is the degree to which deductions from the theory match observations. For example, predictions of growth rates for colonies of yeast.
  • Scope: otherwise referred to as unifying power, is the quality that a theory explains concepts relevant to different phenomena. An example in biology is that we have a unifying theory of genetics for DNA-based life, rather than separate theories for different species or kingdoms.
  • Simplicity: which comes in various forms, depending on the context. Kuhn relates this to tractability using the example of Ptolemy’s and Copernicus’ theory of astronomy and the number of calculations needed for prediction being equivalent in both systems. In some theoretic sense, Copernicus’ theory was simpler than Ptolemy’s, having a simpler mathematical formulation.
  • Fruitfulness (or fertility): which has various interpretations. But can in one way be understood to be: how well does this theory lead to “more science”? When combined with scope we obtain models that generalise well to new applications or other domains. Fruitfulness also has one clear implication for closed-loop discovery.
Importantly, Schindler [30] found that scientists do not agree on the relative importance of these characteristics, although there were some common views. And Kuhn argued that even if there were a common order and weighting, that two individual scientists may honestly differ in their assessment of a better theory because of the ways they evaluate them. This presents a difficulty for robot scientists as well. The difficulty of differences in evaluation is common to human scientists, so we can accept this as part of science. The unique difficulty is imbuing these values into the software used to evaluate theories.
@@
Thankfully, the problem is not as difficult as it may initially appear, as machine learning research is itself informed by similar values. The task then for designing a robot scientist is to align mechanisms from AI research with scientific values in the relevant domain. A domain-specific cost function ensures accuracy is considered; likewise regularisation terms cover simplicity; and there are active research areas in enforcing external consistency on machine learning models, for example imposing symmetry constraints from physics on to neural networks.

### v2

@@ Abstract
The process of developing theories and models, and testing them with experiments is fundamental to the scientific method. Automating the entire scientific method then requires not only automation of the induction of theories from data, but also experimentation from design to implementation. This is the idea behind a robot scientist—a coupled system of AI and laboratory robotics that has agency to test hypotheses with real-world experiments.
In this chapter, we explore some of the fundamentals of robot scientists in the philosophy of science. We also map the activities of a robot scientist to machine learning paradigms, and argue that the scientific method shares an analogy with active learning.
We relate these general guiding principles for designing robot scientists to examples from the domain of system biology: Adam, Eve, and Genesis. We present a case study of Genesis, a next-generation robot scientist designed for research in systems biology, comprising a micro-fluidic system with 1000 computer-controlled micro-bioreactors. We discuss LGEM+, a logic-based model that is used in Genesis, and related them to discussed general principles.
@@ 1 Introduction
In the past two decades, the use of AI-robotic systems in scientific research has been demonstrated not only possible, but fruitful. Several projects—from Adam [20] and Eve [40] to the Robot Chemist [8]—have proved the value of coupling AI software agents with experimental platforms to give them real-world agency. A commonly used term to describe such systems is robot scientist, defined in King et al. [20] as:
“a physically implemented laboratory automation system that exploits techniques from the field of artificial intelligence to execute cycles of scientific experimentation.”
Robot scientists have the potential to broaden and deepen the capability of the scientific community, enabling high-throughput science and new modes of research, as well as helping to address problems with reproducibility, and human resource bottlenecks [41]. Research toward the goal of productive robot scientists is necessarily a multi-disciplinary endeavour. Fields that contribute include: artificial intelligence, robotics, nanotechnology and materials science. For robot scientist projects to be a success, researchers need a shared understanding of the process they are working to automate: the scientific method. This chapter aims to provide researchers with knowledge and tools helpful in analysing and designing robot scientists.
Section 2 first defines some of the core concepts in the philosophical ideas behind robot scientists, beginning with theories and model—the fundamental entities in scientific research. Then, methods of inference are introduced and how these should be considered when designing robot scientists. We explore the concept of parsimony as it relates to the scientific method, something which is learned through scientific education in humans but requires explicit care when creating robot scientists. The section concludes with a description of the different scientific values used to assess the value of a theory or model. Section 3 aims to identify which machine learning paradigm is most apt for scientific discovery. Ultimately, we conclude that scientific discovery shares an analogy with active learning, and that this, rather than reinforcement learning, is the most useful paradigm to adopt when designing or analysing robot scientists. Section 4 gives an overview of systems biology, the domain in which the next generation robot scientist Genesis is applied. We argue that because biological systems are “complex systems” as described by Simon [38], they are excellent targets for study by robot scientists with their superhuman abilities in reasoning and precision. In Section 5 we present a case study of Genesis, and one computational model developed with the aim of automated scientific discovery.
@@ 2 The Philosophy of Robot Scientists
The goal of science is to develop theories that explain and predict phenomena in the real world. To develop theories, science uses models, which are representations of theories in some localised context. Models are a surrogate for the system being studied (object system). This means they have characteristics or behaviour sufficiently similar to that of the object system to allow indirect study of the object system by studying its surrogate. Models are particularly useful when direct study of a system is impossible, impractical or undesirable.
@@
In both of these cases, the object system is the same: a cell. However they have quite different qualities. Model 1 would be well-suited to teaching high school students how cells of the yeast Saccharomyces cerevisiae function. However, Model 2 is capable of quantitative predictions, enabling direct comparison with quantitative experimental data.
@@
Models with deductive capacity using mathematics, such as Model 2, are useful for all forms of scientific discovery, but particularly when automated. The scientific discovery problem becomes, as defined in Flach [11]:
@@
“an incremental process of refinement [of the model] strongly guided by the empirical observations.”
@@
Methods of scientific enquiry rely on: constructing a good starting model; inferring changes to the model; techniques to reason about which models are better; and the collection of relevant, high-quality empirical data to drive the process. Some of these activities are domain-specific, but there are elements that are common among sciences, particularly inference (Section 2.1) and model evaluation (Sections 2.1 and 2.2).
@@
The purpose of a robot scientist is to provide software and hardware that together can achieve each of these activities, and join them together in a “closed-loop” form of enquiry without human intervention [20]. Through the process of designing a system capable of independent scientific inquiry, we seek insight into the scientific method itself, as well as the system subject to enquiry.
2.1 Elements of scientific method
@@
There are three components of the scientific method considered here: logical inference, statistical inference, and parsimony. A history and detailed treatment of each of these components is covered in Gauch [12]; here we provide a brief introduction and highlight how these concepts can be used to analyse robot scientists and how they shape decisions during their design.
2.1.1 Logical inference
Logics are mathematical languages that relate premises and conclusions. Logics are used in science to represent facts (observations of the real world) and laws (parts of theories or models) [9, 12].
There are three basic forms of logical inference: deduction, induction and abduction. With deduction, conclusions are derived from premises and laws. A valid deductive argument is one that guarantees the truth of its conclusions given the truth of its premises. Deduction is what enables deterministic simulations, and how we reason about a hypothesis in the scientific method.
Induction and abduction cannot provide guarantees of truth. Both seek explanations for a set of observable facts. Induction seeks laws that can explain a general case, whereas abduction seeks facts that can explain specific cases.
A simple example of induction is: having observed 10 yeast cultures that thrived in a sugar solution, and another 10 that died in plain water, one induces the law that yeast need sugar to thrive. While this law fits the observations, it turns out to be false in general; some yeasts, including S. cerevisiae can grow on ethanol.
An example of abduction is: given the law induced above, and an observation that a yeast culture is thriving in a liquid of unknown composition, one abduces the fact that there is sugar in the liquid.
Robot scientists require concretely defined induction or abduction problems to make contributions to scientific knowledge. In the case of the first robot scientist, Adam, the scientific problem was to identify links between S. cerevisiae genes and biochemical reactions [39]. By using Enzyme Commission (E.C.) numbers to refer to classes of enzymes which catalyzed a given reaction, the problem was transformed to one of abduction; Adam was to find facts of a given form.
“A single hypothesis is the mapping of one S. cerevisiae ORF [gene] to one E.C. class - e.g. YER152C $\longrightarrow$ 2.6.1.39.”
The discovery algorithm therefore had a well-defined output form. This was possible because the laws applying to E.C. classes were defined in Adam’s logical model. In other cases, when the laws are unknown or not well enough defined, induction will be necessary, and to make these problems concrete one should either use techniques for induction that provide explanations—e.g. inductive logic programming—or ones that do not—e.g. learning neural network models.
Both induction and abduction deal in uncertainty, and therefore need probabilistic and statistical inference.
2.1.2 Statistical inference and probability
Certain laws in science may only be expressed to a degree of certainty, and empirical data are subject to random effects. These situations require statements that can deal with uncertainty, or probability, of which two types are defined by Carnap [9]. Statistical probability is a frequentist idea about the relative frequencies of mass events. The defining characteristic of statements of statistical probability is that they cannot be decided by logic, but rest on empirical observations. Logical probability on the other hand is the probability on a logical relation between two propositions. Gauch prefers to describe the two types of probability as being about events and beliefs respectively [12].
Statistics can be used to obtain a model by reasoning about observations. In the case of abduction, one is reasoning about facts; with induction, laws. The statistical element of this reasoning represents the uncertainty around the model, which can be viewed either as a belief in the law, or perhaps more commonly in empirical science, the relative frequencies of events. Either way, these processes are crucial to the scientific method because they allow us to form and improve upon models.
Statistical reasoning can be done when evaluating hypotheses against empirical data, or during the hypothesis generation stage. Adam, for example, used statistical measures of gene sequence similarity (PSI-BLAST, FASTA) to identify candidate genes in S. cerevisiae from genes in other organisms [39]. And when designing the most recent discovery framework for the robot scientist Eve, Brunnsåker et al. used inductive logic programming to identify candidate phenotypes [7].
@@ 2.1.3 Parsimony
Parsimony is a term that refers to two different concepts. Epistemological parsimony is the concept that when choosing between theories that fit the data equally well, the theory that is simplest is preferable, sometimes referred to as “Ockham’s razor”. Ontological parsimony is the idea that nature itself prefers simplicity. Both are both absolutely essential for science, and it is easy to just take parsimony as a common sense notion. However, a proper treatment of the motivations for adopting parsimony as a guiding principle is warranted when it comes to systems biology and the automation of scientific discovery.
@@
Concrete examples of ontological parsimony include: taking the assumption that there exist common properties between individuals of the same species. This could be a chemical species, that all glucose molecules react with water molecules in the same ways. Or it could be a species of organism, implying that the same computational model can be applied to two individual colonies of S. cerevisiae.
Arguments for ontological parsimony suffer from many counter-examples. For example the yeast genome went through a duplication during its evolution resulting in numerous genes with overlapping or identical function [19]. However all scientific argument must adopt some version of ontological parsimony, as it is the basis for generalisation of theories. Ideas of ontological parsimony are so central to scientific enquiry that researchers may not consider them—they become necessary implicit biases.
@@
One power of appealing to ontological parsimony is that it is the basis of factorial experimentation, described in [9] as a two-step process: firstly, identify relevant factors for the phenomenon to be studied; then design experiments holding certain factors constant and varying over others. Determining relevant factors means stating that certain factors are irrelevant, i.e. that they are not expected to affect the outcome of the experiment. Deciding which factors to include is not an easy process. Designing experiments in this way allows the controlled study of phenomena to test hypotheses relating to a restricted subset of factors. Factorial design of experiments requires an appeal to ontological parsimony, which we can informally express as:
@@
empirical data from experiments of the same class are expected to exhibit only random variation, where the class of experiments depends only on the variable factors.
@@
This assumption allows the inference of empirical laws about the phenomenon.
@@
In practice one cannot usually keep constant all the factors one would like, so the class of experiments does not only depend on relevant variables. The nature of a given experimental protocol will introduce systematic errors: variations in the empirical data not arising from the experimental variables or random noise. Many methods exist to mitigate and model systematic errors, using randomisation techniques, systematic design, and statistics. Randomisation, or systematic designs like Latin squares, can be impractical when using robot scientists due to the limitations of the automation hardware. One example from biology is that liquid handling procedures for Latin square designs are impossible on certain liquid handling robots, and can increase procedure times by orders of magnitude on those with such a capability.
@@
On the other hand, robot scientists have advantages when it comes to relying on and examining Statement (A). Firstly, that robots are capable of performing repeated tasks with a much higher accuracy and precision than human counterparts, as is shown by examples of adoption of laboratory automation in biology [16], physics [33], and chemistry [37]. Secondly, that the validity of Statement (A) can be evaluated by recording more data about the execution of experiments than usually recorded when humans complete experiments. This is a “natural by-product” [20], as robots and software frequently have automatic logging capability. Finally, that those who design robot scientists encode the experimental protocol forces them to specify which experimental factors will be constant, or that if they are not constant they are judged irrelevant, and therefore they do not break the validity of statement above by creating a new class of experiments (for example, the use of different individual glass flasks of the same brand, model, and age).
@@
Epistemological parsimony is a concept much more familiar to human scientists, or at least one which is applied more explicitly. It has been formalised in information theory as the minimum message length (MML): that the hypothesis that best explains the data is that which minimises the total information. The total information has two components, the a priori information contained in the hypothesis, and the information of the data given the hypothesis [^1] Combine Shannon’s formula for the information ($I$) of an event ($E$), given by $I(E)=-log_{2}(P(E))$}, with Bayes’ theorem, $P(E_{1}\cap E_{2})=P(E_{1})P(E_{2}|E_{1})$}. MML states that the hypothesis $H$ that best explains data DD—in other words it maximises $P(H\cap D)$}—is that which minimises the information (message length): $I(H\cap D)=I(H)+I(D|H)$}. . There is a trade-off between these two quantities. One can make a hypothesis so specific as to explain all the data, meaning the information content of the data given the hypothesis is zero. Or one can have no hypothesis at all. In practice, MML ensures that information is only added to a hypothesis if this information explains the data, which is the principle of Ockham’s razor [2, 3, 36].
@@
Epistemological parsimony is the basis for several fundamental concepts in the modern scientific method. One example is the idea of a null hypothesis—a statement that any variance in empirical data can be explained by the extant model, to a degree of certainty due to random noise. To reject the null hypothesis is to accede a more complex model, and is only done if the current model is insufficient to explain the empirical data.
For robot scientists, epistemological parsimony should also be applied in the choice of model as it relates to the experimental hardware. More complex models that suggest hypotheses outside of set of possible experiments result in inaction. Adam was able to achieve autonomous discovery in part because of the parsimony of the logical theory used to generate hypotheses and evaluate them. There were many more facts and relations that could have been included in the theory, but by omitting these and building a theory that was focused on the scientific discovery task, the theory was tractable and resulted only in hypotheses that were testable by the experimental apparatus available to the robot.
@@ 2.2 Comparing scientific models
The relative merit of competing scientific models is not a trivial assessment. This problem is referred to in the philosophy of science as the problem of theory choice. For a robot scientist to be effective, its design must incorporate values and an evaluation procedure, otherwise the scientific discovery process will depend on human evaluation.
@@
  • Internal consistency is defined as the absence of contradictions within a theory. This can be extended to include the various contexts in which the theory is applied; in biology this could mean growth of S. cerevisiae in different conditions.
  • External consistency is the absence of contradictions with other scientific theories. For example, that a model of the biochemistry of yeast is thermodynamically consistent.
  • Empirical accuracy, otherwise referred to as predictive power, is the degree to which deductions from the theory match observations. For example, predictions of growth rates for colonies of yeast.
  • Scope, otherwise referred to as unifying power, is the quality that a theory explains concepts relevant to different phenomena. An example in biology is the existence of a unifying theory of genetics for DNA-based life, rather than separate theories for different species or kingdoms.
  • Simplicity, which comes in various forms, depending on the context. Kuhn relates this to tractability using the example of Ptolemy’s and Copernicus’ theory of astronomy and the number of calculations needed for prediction being equivalent in both systems. In some theoretic sense, Copernicus’ theory was simpler than Ptolemy’s, having a simpler mathematical formulation.
  • Fruitfulness (or fertility), which has various interpretations. But can in one way be understood to be: how well does this theory lead to “more science”? When combined with scope, fruitfulness leads to models that generalise well to new applications or other domains. Fruitfulness also has one clear implication for closed-loop discovery.
Schindler [35] found that scientists do not agree on the relative importance of these characteristics, although there were some common views. And Kuhn [23] argued that even if there were a common order and weighting, that two individual scientists may honestly differ in their assessment of a better theory because of the ways they evaluate them. This presents a difficulty for robot scientists as well. The difficulty of differences in evaluation is common to humans, so we can accept this as part of science. The unique difficulty for robot scientists is imbuing these values into the software used to evaluate theories.
@@
Thankfully, the problem is not as difficult as it may initially appear, as machine learning research is itself informed by similar values. The task then for designing a robot scientist is to align mechanisms from AI research with scientific values in the relevant domain. A domain-specific cost function ensures accuracy is considered; likewise regularisation terms cover simplicity; and there are active research areas in enforcing external consistency on machine learning models, for example imposing symmetry constraints from physics on to neural networks[1].